OpenStudy (anonymous):

lim x-> pi/4 of (1-tanx)/(sinx-cosx)

OpenStudy (blockcolder):

Are you allowed to use L'hospital's Rule? If so, use it.

OpenStudy (anonymous):

i have no idea what that is

OpenStudy (blockcolder):

Then I guess you aren't. Try multiplying both numerator and denominator by cos(x).

OpenStudy (anonymous):

mess around with the expression until you turn it in to \(-\frac{1}{\sin(x)}\)

OpenStudy (anonymous):

okay i know tan x = sins/cos x

OpenStudy (anonymous):

so 1- (sinx/ cos x)/ sinx- cos x